# The role of individual compensation and acceptance decisions in crowdsourced delivery

### v4

@@ 1 Introduction
Crowdsourced delivery adds significant complexity to the planning process because the behavior and availability of occasional drivers is often not known in advance. Integrating the behavior of occasional drivers into planning decisions, especially with respect to task acceptance, is a major challenge in this domain (Savelsbergh and Ulmer, 2022). In the case of instant delivery (e.g., food orders), the inability to find an occasional driver for an offered task may lead to customer dissatisfaction or significant additional costs. In the classical models, occasional drivers are assumed to accept all tasks for which the compensation is high enough to compensate for the additional costs incurred by the drivers for the detour, see e.g. Archetti et al. (2016); Arslan et al. (2019). More recently, acceptance probabilities have been investigated through behavioral studies and surveys of occasional drivers, see e.g. Devari et al. (2017); Le and Ukkusuri (2019). Such acceptance probabilities were used as fixed parameters in the task allocation problem by Gdowska et al. (2018) and Santini et al. (2022). The effect of compensation schemes on acceptance behavior was investigated by Dayarian and Savelsbergh (2020), but only through a sensitivity analysis of fixed compensation options and acceptance thresholds drawn from a normal distribution. Barbosa et al. (2022) and Hou et al. (2022) integrate the impact of compensation on acceptance probabilities within task allocation. However, Barbosa et al. (2022) assume that the detour of all occasional drivers is the same for each task. Consequently, the compensations of all occasional drivers are identical. Hou et al. (2022) propose a sequential approach in which the compensation values are optimized for fixed task allocations that are decided in the first step.
@@ 5 The CAPCAB without separability
Proposition 4 implies that a piecewise linear approximation of the nonlinear objective function (2a) can be derived using standard techniques as described, e.g., in Nemhauser and Wolsey (1988). For each task $i\in I$ and every driver $j\in J$, we approximate the non-linear function $f_{ij}(C_{ij})$ for $C_{ij}\in[0,U_{ij}]$ by a piecewise linear function defined using a discrete set of $K$ possible compensation values $u_{ij}^{k}$, $k=1,\dots,K$. Formally, the approximation is specified by the points $(u_{ij}^{k},f_{ij}(u_{ij}^{k}))$, $k=1,\dots,K$, where $u_{ij}^{1}=0$, $u_{ij}^{K}=U_{ij}$, and $u_{ij}^{k-1}<u_{ij}^{k}$, $k\in\{2,\dots,K\}$. For all $i\in I,j\in J$ and $k\in\{1,\dots,K\}$ we further define non-negative auxiliary variables $\lambda_{ij}^{k}\geq 0$ as the weights for the convex combination of the values of $f_{ij}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C}_{ij})$. Therefore we replace $f_{ij}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C}_{ij})$ in (13) by the convex combination $\sum_{k=1}^{K}\lambda_{ij}^{k}f_{ij}^{k}(u_{ij}^{k})$ such that $\sum_{k=1}^{K}\lambda_{ij}^{k}=1$. Since the nonlinear functions $f_{ij}(C_{ij})$ are not necessarily convex, we need to additionally ensure that at most two of these variables are positive and that two such variables can only be both positive if they are consecutive, i.e., if $\lambda_{ij}^{k}>0$ and $\lambda_{ij}^{r}>0$, then $r\in\{k-1,k+1\}$. To ensure these additional conditions, formulation (14) also uses binary variables $v_{ij}^{k}\in\{0,1\}$ for each $i\in I$, $j\in J$, and $k\in\{1,\dots,K-1\}$. Here, $v_{ij}^{k}=1$ indicates that $u_{ij}^{k}\leq C_{ij}\leq u_{ij}^{k+1}$.
$\min\quad$ $\sum_{i\in I}c_{i}y_{i}+\sum_{i\in I}\sum_{j\in J}\left(\sum_{k=1}^{K}f_{ij}(u_{ij}^{k}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{ij}^{k}}+g_{ij}(x_{ij})\right)$ (14a)
@@
$\sum_{k=1}^{K}u_{ij}^{k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{ij}^{k}}\leq U_{ij}x_{ij}$ $i\in I,\ j\in J$ (14b)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{ij}^{1}}\leq v_{ij}^{1}$ $i\in I,\ j\in J$ (14c)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{ij}^{k}}\leq v_{ij}^{k-1}+v_{ij}^{k}$ $i\in I,j\in J,\ k\in\{2,\dots,K-1\}$ (14d)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{ij}^{K}}\leq v_{ij}^{K-1}$ $i\in I,\ j\in J$ (14e)
@@
$\sum_{k=1}^{K}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda_{ij}^{k}}=1$ $i\in I,\ j\in J$ (14g)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda}_{ij}^{k}\geq 0$ $i\in I,\ j\in J,\ k\in\{1,\dots,K\}$ (14i)
The objective function (14a) is a piecewise linear approximation of the one introduced in Proposition 4 using the concepts described above. Recall that $g_{ij}(x_{ij})$ is a linear function. As described in Section 3.1, constraints (2b)-(2f) ensure that at most one task is offered to an occasional driver and that each task is either performed by a professional driver, or offered to an occasional driver. They also define the domains of variables $x_{ij}$ and $y_{i}$ for tasks $i\in I$ and occasional drivers $j\in J$. Constraints (12) have been introduced at the beginning of this section while inequalities (14b) are constraints $C_{ij}\leq U_{ij}x_{ij}$ rewritten using the identity $\sum_{k=1}^{K}u_{ij}^{k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda}_{ij}^{k}=C_{ij}$. Finally, (14c)–(14i) are standard constraints used to model piecewise linear approximations that ensure the conditions detailed above, see, e.g., Nemhauser and Wolsey (1988) for further details.

### v5

@@ 1 Introduction
Crowdsourced delivery adds significant complexity to the planning process because the behavior and availability of occasional drivers is often not known in advance. Integrating the behavior of occasional drivers into planning decisions, especially with respect to task acceptance, is a major challenge in this domain (Savelsbergh and Ulmer, 2022). In the case of instant delivery (e.g., food orders), the inability to find an occasional driver for an offered task may lead to customer dissatisfaction or significant additional costs. In the classical models, occasional drivers are assumed to accept all tasks for which the compensation is high enough to compensate for the additional costs incurred by the drivers for the detour, see e.g. Archetti et al. (2016); Arslan et al. (2019). More recently, acceptance probabilities have been investigated through behavioral studies and surveys of occasional drivers, see e.g. Devari et al. (2017); Le and Ukkusuri (2019). Such acceptance probabilities were used as fixed parameters in the task allocation problem by Gdowska et al. (2018) and Santini et al. (2022). The effect of compensation schemes on acceptance behavior was investigated by Dayarian and Savelsbergh (2020), but only through a sensitivity analysis of fixed compensation options and acceptance thresholds drawn from a normal distribution. Le et al. (2021) propose a compensation scheme that defines occasional drivers’ behavior as threshold constraints. Thresholds are identified through survey data analysis and any offer above these thresholds is assumed to be accepted. Barbosa et al. (2022) and Hou et al. (2022) integrate the impact of compensation on acceptance probabilities within task allocation. However, Barbosa et al. (2022) assume that the detour of all occasional drivers is the same for each task. Consequently, the compensations of all occasional drivers are identical. Hou et al. (2022) propose a sequential approach in which the compensation values are optimized for fixed task allocations that are decided in the first step. Our study differs from Le et al. (2021) by using probability functions to describe occasional driver behavior instead of threshold constraints; from Barbosa et al. (2022) by considering the characteristics of each occasional driver-task pair rather than assuming identical values for all; and from Hou et al. (2022) by evaluating all possible occasional driver task combinations and compensation values to find a global optimal solution, rather than using a sequential approach without a guarantee of optimality.
@@ 3.3 Logistic acceptance probability
3.4 Learning acceptance probabilities from historical data
Crowdsourced delivery platforms use online tools to process delivery orders and allocate them to occasional drivers. This enables them to collect a wide range of data about their operations. In the following, we discuss how historical data about offers made to occasional drivers can be used to learn acceptance probabilities that can be used within our framework. For this, the data collected for previous offers must contain information about (i) the task, (ii) the driver, (iii) the compensation offered to the driver, and (iv) whether the driver accepted the offer as illustrated in Table 2.
The required information associated with tasks and drivers depends on the operational setting and considered probability function. For the in-store setting described in the previous sections, it is sufficient to keep track of task delivery points and driver destinations to calculate the detour, since both, task shipment points of the tasks as well as driver initial locations are assumed to be the store. In a setting where tasks have distinct shipment points and drivers have distinct initial locations, keeping track of such additional information is required to calculate the detour. As described below, additional information about the environment must be included in the database in case their impact should be incorporated in the acceptance probability functions.
Table 2: A snapshot of representative operational data
Task information (i) Driver information (ii) Offer information
ID Shipment point Delivery point ID Location Destination Compensation (iii) Detour Response (iv)
T001 P001 P003 D001 P005 P006 3.7 EUR 5.3 KM Reject
T002 P002 P004 D002 P002 P007 5.9 EUR 1.2 KM Accept
T003 P001 P004 D003 P001 P008 4.1 EUR 3.8 KM Reject
A logistic (or linear) acceptance probability model can be developed using such historical data to estimate driver responses based on task and driver information, as well as offered compensation as independent variables. In such a model, the impact of compensation is captured by the $\delta$ (or $\beta$) parameter, while the combined influence of other factors is represented by the $\gamma$ (or $\alpha$) parameter in Equation 5 (or Equation 3). This approach also allows to easily incorporate additional factors related to drivers, orders, or the environment using the $\gamma$ ($\alpha$) parameter. Considering additional factors is especially important for revealing the indirect relationships that impact driver behavior. While the detour could be a significant factor that impact drivers’ decisions during non-rush hours, the expected delivery time could, in contrast, be more significant during rush hours. Similarly, weather conditions can impact driver response depending on the vehicle used by the driver. A driver using bikes to make deliveries could be less willing to make extra deliveries on a rainy day. Hence, integrating such factors can improve predicting drivers’ responses. In practice, such parameters can be easily integrated into the proposed methodology without changing the formulations.
Fitting accurate acceptance probability models at the individual level may be challenging due to potentially insufficient data for each individual. Instead, one can identify behavioral clusters among drivers and fit acceptance probability models that describe the behavior of each cluster. Mohri et al. (2024) demonstrates such an approach by employing a latent class model. Alternatively, one can use the complete dataset to fit an acceptance probability model that captures the effects of the observable factors by including them as variables in the model. Such an approach is demonstrated by Le and Ukkusuri (2019) to evaluate the effects of various socioeconomic factors on drivers’ decisions to participate in crowdsourced delivery.
@@ 5 The CAPCAB without separability
Proposition 4 implies that a piecewise linear approximation of the nonlinear objective function (2a) can be derived using standard techniques as described, e.g., in Nemhauser and Wolsey (1988). For each task $i\in I$ and every driver $j\in J$, we approximate the non-linear function $f_{ij}(C_{ij})$ for $C_{ij}\in[0,U_{ij}]$ by a piecewise linear function defined using a discrete set of $K$ possible compensation values $u_{ij}^{k}$, $k=1,\dots,K$. Formally, the approximation is specified by the points $(u_{ij}^{k},f_{ij}(u_{ij}^{k}))$, $k=1,\dots,K$, where $u_{ij}^{1}=0$, $u_{ij}^{K}=U_{ij}$, and $u_{ij}^{k-1}<u_{ij}^{k}$, $k\in\{2,\dots,K\}$. For all $i\in I,j\in J$ and $k\in\{1,\dots,K\}$ we further define non-negative auxiliary variables $\lambda_{ij}^{k}\geq 0$ as the weights for the convex combination of the values of $f_{ij}({C}_{ij})$. Therefore we replace $f_{ij}({C}_{ij})$ in (13) by the convex combination $\sum_{k=1}^{K}\lambda_{ij}^{k}f_{ij}^{k}(u_{ij}^{k})$ such that $\sum_{k=1}^{K}\lambda_{ij}^{k}=1$. Since the nonlinear functions $f_{ij}(C_{ij})$ are not necessarily convex, we need to additionally ensure that at most two of these variables are positive and that two such variables can only be both positive if they are consecutive, i.e., if $\lambda_{ij}^{k}>0$ and $\lambda_{ij}^{r}>0$, then $r\in\{k-1,k+1\}$. To ensure these additional conditions, formulation (14) also uses binary variables $v_{ij}^{k}\in\{0,1\}$ for each $i\in I$, $j\in J$, and $k\in\{1,\dots,K-1\}$. Here, $v_{ij}^{k}=1$ indicates that $u_{ij}^{k}\leq C_{ij}\leq u_{ij}^{k+1}$.
$\min\quad$ $\sum_{i\in I}c_{i}y_{i}+\sum_{i\in I}\sum_{j\in J}\left(\sum_{k=1}^{K}f_{ij}(u_{ij}^{k}){\lambda_{ij}^{k}}+g_{ij}(x_{ij})\right)$ (14a)
@@
$\sum_{k=1}^{K}u_{ij}^{k}{\lambda_{ij}^{k}}\leq U_{ij}x_{ij}$ $i\in I,\ j\in J$ (14b)
${\lambda_{ij}^{1}}\leq v_{ij}^{1}$ $i\in I,\ j\in J$ (14c)
${\lambda_{ij}^{k}}\leq v_{ij}^{k-1}+v_{ij}^{k}$ $i\in I,j\in J,\ k\in\{2,\dots,K-1\}$ (14d)
${\lambda_{ij}^{K}}\leq v_{ij}^{K-1}$ $i\in I,\ j\in J$ (14e)
@@
$\sum_{k=1}^{K}{\lambda_{ij}^{k}}=1$ $i\in I,\ j\in J$ (14g)
@@
${\lambda}_{ij}^{k}\geq 0$ $i\in I,\ j\in J,\ k\in\{1,\dots,K\}$ (14i)
The objective function (14a) is a piecewise linear approximation of the one introduced in Proposition 4 using the concepts described above. Recall that $g_{ij}(x_{ij})$ is a linear function. As described in Section 3.1, constraints (2b)-(2f) ensure that at most one task is offered to an occasional driver and that each task is either performed by a professional driver, or offered to an occasional driver. They also define the domains of variables $x_{ij}$ and $y_{i}$ for tasks $i\in I$ and occasional drivers $j\in J$. Constraints (12) have been introduced at the beginning of this section while inequalities (14b) are constraints $C_{ij}\leq U_{ij}x_{ij}$ rewritten using the identity $\sum_{k=1}^{K}u_{ij}^{k}{\lambda}_{ij}^{k}=C_{ij}$. Finally, (14c)–(14i) are standard constraints used to model piecewise linear approximations that ensure the conditions detailed above, see, e.g., Nemhauser and Wolsey (1988) for further details.